# A polynomially accelerated fixed-point iteration for vector problems

### v1

@@ Abstract
Abstract  Fixed-point procedures frequently slow down because an error mode decays much more slowly than the others, leaving the base iteration with a persistent residual plateau. To counter this obstruction we formulate a three-point polynomial accelerator (TPA) that fits inside existing fixed-point algorithms with negligible modification and computational cost. TPA first infers the dominant contraction factor directly from the residual dynamics and then assembles a regularised three-point update from the last three iterates. We show that across a suite of tests: a linear system with clustered eigenvalues, a nonlinear $\tanh$ mapping, and a discretised Poisson equation, TPA attains a prescribed tolerance in markedly fewer map evaluations than Picard iteration, weighted Jacobi/SOR, and shallow Anderson schemes while preserving a minimal memory and arithmetic footprint.
@@ 1 Introduction
From a formal standpoint, the analysis that leads to TPA builds upon the classical literature on sequence transformations. Aitken’s $\Delta^{2}$ process removes the leading error term in one dimension [aitken_xxv.bernoullis_1927]; minimal-polynomial extrapolation transfers the annihilating-polynomial perspective to vector-valued settings [sidi_vector_2008]; and Anderson acceleration constructs optimal mixing coefficients via constrained least squares [anderson_iterative_1965]. Successive over-relaxation [young_iterative_1954] remains the canonical spectral shaper but hinges on careful damping-parameter selection. A comprehensive synthesis of these acceleration paradigms and their extensions is presented in the survey [saad_acceleration_2025]. In order to make this bridge between our algorithm and pre-existing methods, in Sections 4.1 and 4.2 we summarise its precise relation to Aitken’s $\Delta^{2}$ process and Anderson acceleration of depth two.
@@ 2 Derivation of the Algorithm
When $T$ is differentiable at $x_{\star}$, let $J=T^{\prime}(x_{\star})$ denote its Jacobian. The linearisation gives $r_{n+1}=Jr_{n}+o(\|e_{n}\|)$. If $J$ admits a dominant mode with right and left vectors $u$ and $v$, then
@@ 3.1 Linear System with Clustered Spectrum
TPA 32 $0.16\times 10^{8}$ $0.12\times 10^{6}$
Picard 1197 $1.00\times 10^{8}$ $0.98\times 10^{6}$
SOR ($\omega=1.80$) 663 $0.99\times 10^{8}$ $0.97\times 10^{6}$
Anderson(m=2) 419 $0.98\times 10^{8}$ $0.86\times 10^{6}$
Anderson(m=3) 183 $0.95\times 10^{8}$ $0.94\times 10^{6}$
Anderson(m=5) 52 $0.94\times 10^{8}$ $0.72\times 10^{6}$
@@ 3.2 Nonlinear tanh Fixed Point
TPA 36 $0.53\times 10^{8}$ $0.18\times 10^{7}$
Picard 128 $0.97\times 10^{8}$ $0.91\times 10^{7}$
SOR ($\omega=1.50$) 84 $0.85\times 10^{8}$ $0.79\times 10^{7}$
Anderson(m=2) 70 $0.86\times 10^{8}$ $0.64\times 10^{7}$
Anderson(m=3) 49 $0.89\times 10^{8}$ $0.72\times 10^{7}$
Anderson(m=5) 38 $0.69\times 10^{8}$ $0.49\times 10^{7}$
@@ 3.3 Discretised 2D Poisson Equation
Table 3: Discretised 2D Poisson equation. Performance on the $50\times 50$ grid for Jacobi, weighted Jacobi, Anderson, and TPA.
@@
TPA 244 $0.97\times 10^{8}$ $0.51\times 10^{5}$
Picard 4317 $1.00\times 10^{8}$ $0.53\times 10^{5}$
Anderson(m=2) 3722 $1.00\times 10^{8}$ $0.51\times 10^{5}$
Anderson(m=3) 1573 $0.99\times 10^{8}$ $0.47\times 10^{5}$
Anderson(m=5) 955 $0.96\times 10^{8}$ $0.24\times 10^{5}$

### v2

@@ Abstract
Abstract   Fixed-point solvers are ubiquitous in nonlinear PDEs, yet their progress collapses whenever the Jacobian at the solution carries an eigenvalue arbitrarily close to one. We ask whether such stagnation can be removed without storing long histories or solving dense least squares. Under two assumptions—(A1) the linearised error $e_{n}$ is dominated by a multiplier $m$ with $|m|<1$ and (A2) residuals shrink monotonically—we construct a quadratic blend of three iterates whose error polynomial has a double root at $m$. This three-point polynomial accelerator (TPA) cancels the stubborn mode up to $o(\|e_{n}\|)$, reduces to Aitken’s $\Delta^{2}$ process in one dimension, and matches a doubly blended Anderson step with depth $m=2$ when the regularisation vanishes, yet it keeps the Picard memory footprint. The only extra ingredient is a residual-based estimate of $w=(1-m)^{-1}$ obtained from a closed-form regularised least-squares fit that remains stable even when two residuals nearly coincide. Numerical experiments on linear systems with clustered spectra, a $320$-dimensional nonlinear $\tanh$ fixed point, and a $50\times 50$ Poisson discretisation show that TPA reaches the $10^{-8}$ residual tolerance in $32$, $36$, and $244$ map evaluations (respectively). In the same settings SOR requires $663$ steps and Anderson acceleration with depth $m=5$ consumes $52$, $38$, and $955$ evaluations. TPA therefore supplies a parameter-free, constant-memory drop-in accelerator whenever a single contraction factor throttles convergence.
Keywords: fixed-point iteration; acceleration methods; three-point polynomial accelerator; Anderson acceleration; polynomial extrapolation.
MSC 2020: 65H10; 65F10; 65B05.
@@ 1 Introduction
From a formal standpoint, the analysis that leads to TPA builds upon the classical literature on sequence transformations. Aitken’s $\Delta^{2}$ process removes the leading error term in one dimension [1]; minimal-polynomial extrapolation transfers the annihilating-polynomial perspective to vector-valued settings [4]; and Anderson acceleration constructs optimal mixing coefficients via constrained least squares [2]. Successive over-relaxation [5] remains the canonical spectral shaper but hinges on careful damping-parameter selection. A comprehensive synthesis of these acceleration paradigms and their extensions is presented in the survey [3]. In order to make this bridge between our algorithm and pre-existing methods, in Sections 4.1 and 4.2 we summarise its precise relation to Aitken’s $\Delta^{2}$ process and Anderson acceleration of depth two.
@@ 2 Derivation of the Algorithm
When $T$ is differentiable at $x_{\star}$, let $J=T^{\prime}(x_{\star})$ denote its Jacobian. The linearisation gives $e_{n+1}=Je_{n}+o(\|e_{n}\|)$. If $J$ admits a dominant mode with right and left vectors $u$ and $v$, then
@@ 3.1 Linear System with Clustered Spectrum
TPA 32 $0.16\times 10^{-8}$ $0.12\times 10^{-6}$
Picard 1197 $1.00\times 10^{-8}$ $0.98\times 10^{-6}$
SOR ($\omega=1.80$) 663 $0.99\times 10^{-8}$ $0.97\times 10^{-6}$
Anderson(m=2) 419 $0.98\times 10^{-8}$ $0.86\times 10^{-6}$
Anderson(m=3) 183 $0.95\times 10^{-8}$ $0.94\times 10^{-6}$
Anderson(m=5) 52 $0.94\times 10^{-8}$ $0.72\times 10^{-6}$
@@ 3.2 Nonlinear tanh Fixed Point
TPA 36 $0.53\times 10^{-8}$ $0.18\times 10^{-7}$
Picard 128 $0.97\times 10^{-8}$ $0.91\times 10^{-7}$
SOR ($\omega=1.50$) 84 $0.85\times 10^{-8}$ $0.79\times 10^{-7}$
Anderson(m=2) 70 $0.86\times 10^{-8}$ $0.64\times 10^{-7}$
Anderson(m=3) 49 $0.89\times 10^{-8}$ $0.72\times 10^{-7}$
Anderson(m=5) 38 $0.69\times 10^{-8}$ $0.49\times 10^{-7}$
@@ 3.3 Discretised 2D Poisson Equation
Table 3: Discretised 2D Poisson equation. Performance on the $50\times 50$ grid for Jacobi (Picard), Anderson, and TPA.
@@
TPA 244 $0.97\times 10^{-8}$ $0.51\times 10^{-5}$
Picard 4317 $1.00\times 10^{-8}$ $0.53\times 10^{-5}$
Anderson(m=2) 3722 $1.00\times 10^{-8}$ $0.51\times 10^{-5}$
Anderson(m=3) 1573 $0.99\times 10^{-8}$ $0.47\times 10^{-5}$
Anderson(m=5) 955 $0.96\times 10^{-8}$ $0.24\times 10^{-5}$